factor and solve
3x^2-11x=20

Answer by helpnalgebra(91)

3x^(2)-11x=20
3x^(2)-11x-20=0
(x+(4)/(3))(x-5)=0
(3x+4)(x-5)=0
3x+4=0
3x=-4
3x/3=-4/3
x=-4/3
x-5=0
x=5
x=-4/3,5